OpenStudy (maxdrizner):

Which of the following is an equivalent form of the compound inequality -44 > -2x - 8 ≥ -8

OpenStudy (anonymous):

-44 > -2x - 8 -44 + 8 > -2x - 36 > -2x -36/-2 < x (change the sign when dividing by a negative) 18 < x -2x - 8 >= -8 8 - 8 >= 2x 0 >= 2x 0/2 >= x 0 >= x so we have x is greater then 18 and x is less then or equal to 0 open circle on 18, shading to the right and closed circle on 0, shading to the left.

OpenStudy (anonymous):

any questions ?